# Stability and dynamics of massive vortices in two-component Bose-Einstein condensates

### v1

@@ I Introduction
The platform of atomic Bose-Einstein condensates (BECs) has been particularly fruitful towards the exploration of nonlinear wave patterns, both solitonic ones, as well as vortical ones involving topological charge Pethick and Smith (2002); Stringari and Pitaevskii (2003); Pitaevskii and Stringari (2003). The ability to manipulate external potentials in both space and time Henderson et al. (2009), as well as that of controlling the size and the sign of the atomic scattering length (and hence the coefficient of nonlinear interaction) in time Donley et al. (2001); Pollack et al. (2009) or even in space Kartashov et al. (2011); Di Carli et al. (2020) have paved new directions of unprecedented perturbations and dynamics to the solitary wave patterns that such systems can support.
@@
While these higher-component settings offer considerable additional wealth in terms of the possible states, our aim herein is to revisit two-component settings. There, many solitary waveforms have been considered, ranging from dark-bright (DB) to dark-dark and dark-antidark, as well as magnetic solitons Kevrekidis et al. (2015); Kevrekidis and Frantzeskakis (2016); Qu et al. (2016). Not only individual structures of this kind have been experimentally explored, but also small clusters thereof (see, e.g., Ref. Katsimiga et al. (2020)), as well as their collisions Becker et al. (2008). Such two-component settings have recently been leveraged also towards engineering effective attractive interactions in both 1D and 2D, offering the possibility to generate structures such as the Townes soliton Bakkali-Hassani et al. (2021), as well as the Peregrine soliton Romero-Ros et al. (2024).
Here, we focus on the 2D variants of the above solitary waves in the form of vortex-bright (VB) solitons Skryabin (2000); Law et al. (2010). These are sometimes referred to by other names such as baby-skyrmions, filled-core vortices, or massive vortices and have not only been theoretically explored but also experimentally observed Anderson et al. (2000); Leanhardt et al. (2003); Choi et al. (2012). Note that some of these experimental realizations such as the one reported in Ref. Choi et al. (2012) (and corresponding theoretical explorations such as in Ref. Katsimiga et al. (2023)) have probed these states in spinor settings. In recent years there has been a flurry of activity in such systems, seeking to provide a theoretical understanding of the potential dynamics of these structures. Indeed, they have been shown to be able to perform flower-like motions Richaud et al. (2021); Caldara et al. (2023); they have been also explored in generalized radial potentials Wang (2022); Richaud et al. (2022), as well as in the context of rotating traps Richaud et al. (2020) and more recently the case of multiple such states and how they interact has been revisited Richaud et al. (2023) (see also earlier studies in Refs. Pshenichnyuk (2018); Kasamatsu et al. (2016)).
@@ II.1 Full PDE model
The VB states are studied in the parameter space. Here, the key parameters determining the properties of the VB complex are $\Omega$, $g_{12}$, $\mu_{d}$, and $\mu_{b}$. For example, the powers and the mass ratio depend directly, in a non-trivial manner, on the chemical potentials for a given set of $\Omega$ and $g_{12}$ values.
We find numerically exact stationary states using finite differencing for the spatial discretization together with a Newton-based fixed-point iteration method. First, we find a relevant solution, and then numerical continuation can be deployed along any of the above parameters. The first solution is obtained by embedding a vortex centered at the origin in the first component, we shall discuss the details in Sec. III. We also study the (spectral) stability of these centered VB complexes by the BdG analysis. Then, taking parameter values where the VB complex is stable, we proceed to appropriately seed it away from the trap center and monitor, according to Eq. (1), its dynamical evolution. Our dynamics is carried out using the regular fourth-order Runge-Kutta method. The location of the VB complex is monitored by computing the respective centers of mass of each component along the $x$ and $y$ directions:
@@
where $\sigma=x,y$. Then, the relative location of the VB with respect to the background cloud of the dark component is recorded.
@@
$-i\left[\begin{array}[]{cccc}\phantom{-}M_{11}&\phantom{-}M_{12}&\phantom{-}M_{13}&\phantom{-}M_{14}\\ -M_{12}^{*}&-M_{11}^{*}&-M_{14}^{*}&-M_{13}^{*}\\ \phantom{-}M_{13}^{*}&\phantom{-}M_{14}&\phantom{-}M_{22}&\phantom{-}M_{34}\\ -M_{14}^{*}&-M_{13}&-M_{34}^{*}&-M_{22}^{*}\\ \end{array}\right]\left[\begin{array}[]{c}a_{1}\\ b_{1}\\ a_{2}\\ b_{2}\end{array}\right]=\nu\left[\begin{array}[]{c}a_{1}\\ b_{1}\\ a_{2}\\ b_{2}\end{array}\right],$ (7)
@@
and $M_{12}\equiv\psi_{1}^{2}$, $M_{13}\equiv g_{12}\psi_{1}\psi_{2}^{*}$, $M_{14}\equiv g_{12}\psi_{1}\psi_{2}$, and $M_{34}\equiv\psi_{2}^{2}$. The eigenvector for the corresponding eigenvalue $\nu$ is $[\Psi_{1},\Psi_{2}]^{T}=[(a_{1},b_{1}),(a_{2},b_{2})]^{T}$, and the mode is unstable if Re$(\nu)>0$. As we will see in Sec. III, after stationary solutions are displaced away from the origin and for appropriate parameter regimes, the VB states can exhibit flowering trajectories whose two main frequencies are approximated by the pertinent eigenvalues of the corresponding BdG spectrum. In order to follow potential instabilities emerging from the collision of eigenvalues, for each eigenvector we monitor the sign of the quantity (so-called Krein signature) Kapitula et al. (2004):
@@ II.2 Reduced ODE model for the VB complex
Note that we may also write $L$ in polar coordinates using standard conversion formulas $x^{2}+y^{2}=r^{2}$, $\dot{x}^{2}+\dot{y}^{2}=\dot{r}^{2}+r^{2}\dot{\theta}^{2}$, and $x\dot{y}-y\dot{x}=r^{2}\dot{\theta}$ for $r=r(t)$ and $\theta=\theta(t)$. Comparing Eq. (11) above with the Lagrangian (23) of Ref. Richaud et al. (2021), we find similar structural features with the expression used herein bearing an additional contribution $\left[\mu_{d}-\frac{1}{2}V(r)\right]$ in the second kinetic energy term, while instead of their logarithm, we include herein only the quadratic term in its Taylor series. Additionally, their $r$ is normalized as $r/R_{\rm TF}$ for $R_{\rm TF}$ the TF radius of the dark component.
@@
On the other hand, to derive the Hamiltonian formulation of the ODE system (10) we can utilize the Legendre transform $H=\dot{x}\rho_{x}+\dot{y}\rho_{y}-L$ with $\rho_{\sigma}=\frac{\partial L}{\partial\dot{\sigma}}$ for $\sigma=x,y$, to obtain
$H$ $=$ $\frac{1}{2M_{0}}\left(\rho_{x}^{2}+\rho_{y}^{2}-(y\rho_{x}-x\rho_{y})\left(\mu_{d}-\frac{1}{2}V(r)\right)\right)$ (13)
$+\frac{r^{2}}{8M_{0}}\left(\mu_{d}-\frac{1}{2}V(r)\right)-\frac{\Lambda}{4}r^{2}.$
@@
$V_{\rm eff}(r)=\frac{1}{2}M_{0}r^{2}\dot{\theta}^{2}-\frac{\Lambda}{4}r^{2}.$ (15)
@@
$M_{0}(\ddot{r}-r\dot{\theta}^{2})=-\rho(r)r\dot{\theta}+\frac{\Lambda}{2}r,$ (17)
@@
The smaller frequency $\omega_{\rm pr}\equiv\dot{\theta}_{-}$ corresponds to the expected precessional frequency of a circle, but we also obtain the larger frequency $\dot{\theta}_{+}$ that corresponds to the radial or flowering frequency.
@@
$\dot{y}(0)=\frac{1}{2M_{0}}\left(\rho(r_{0})x_{0}\pm\sqrt{\rho(r_{0})^{2}x_{0}^{2}-2M_{0}\Lambda r_{0}^{2}}\right).$ (20)
@@
for radial and precesional frequencies $(\omega_{r},\omega_{\rm pr})$ and $R$ and $A_{0}$ being small perturbations off of the circular motion. Plugging Eqs. (LABEL:eq:estRTheta) into the first order system above gives
@@
In the next section, we will make the correspondence between the PDE model and its effective ODE reduction using the following program. Since we expect frequencies $\dot{\theta}$ from the ODE model to correspond to purely imaginary frequencies in the PDE model, we use the BdG bifurcating eigenvalues $\nu$ and the correspondence $\dot{\theta}={\rm Im}(\nu)$ in order to estimate the ODE parameter value of $\Lambda$ in terms of $M_{0}$. The value of $M_{0}$ is then determined by minimizing the sum of squares of differences between the PDE and ODE trajectories. For a given pair of $M_{0}$ and $\Lambda$ one then finds initial velocities of the ODE model that best align to a given PDE trajectory. In general, a larger initial velocity gives flowering trajectories with more looping on the petals so this can be used to guide best choices of these initial velocities in the ODE. For the obtained flowering trajectories, we will also show examples to compare the ODE trajectories according to Eq. (10) versus the results of the linearized predictions in Eqs. (LABEL:eq:estRTheta) and (LABEL:eq:omegaA).
@@ III.1 Steady states and stability for the PDE model
Let us now track the BdG eigenvalue spectrum for the centered VB structure as a function of the parameter $\mu_{b}$. Figure 5 depicts a typical spectrum for $g_{12}=1.6$ and $\mu_{d}=4.25$. Note that Re$(\nu)>0$ indicates an instability. Interestingly, as the spectrum shows, there is a Hamiltonian Hopf bifurcation as $\mu_{b}$ decreases. For $\mu_{d}=4.25$ and values $g_{12}=1.4,1.6,1.75$, the bifurcations occur, respectively, at $\mu_{b}\approx 4.349,4.383,4.403$. This (reverse) Hamiltonian Hopf bifurcation evidences that the lowest two oscillatory (purely imaginary) eigenfrequencies collide as $\mu_{b}$ decreases giving rise to an unstable (complex) quartet. The presence of a complex quartet instability indicates that, in this parametric regime, the VB will destabilize from the center of the trap along an spiralling-out trajectory as we will see below soon. According to the color-coded eigenvalues in Fig. 5, the region of instability lies where the lower branch is colored black for $\mu_{b}\lesssim 4.383$. However, it is interesting that, with the parameters provided in Fig. 5, spiralling-out is also present for $\mu_{b}$ values as high as $\mu_{b}<4.4$ when the VB structure is displaced by a large value such as $x_{0}=2.5$. This tends to suggest that precessional circular orbits may also be prone to instabilities. This is a topic of particular interest in its own right, meriting an examination of such states in a co-rotating frame (as was done, e.g., in the work of Pelinovsky and Kevrekidis (2012)), yet this type of distinct study is an interesting topic for future work.
@@
Focusing on initially centered VB solutions, according to our BdG spectral analysis, starting infinitesimally close to the center, spiral-out trajectories will only be present in the $\mu_{b}$-interval where the stationary solutions are unstable. For instance, the VB dynamics for an initial condition $(x_{0},y_{0})=(2.5,0)$ in the stable regime ($\mu_{b}\approx 4.4$) displays flowering trajectories that have roughly two petals and as $\mu_{b}$ increases the number of petals increase. In Fig. 5, we overlay the two frequencies obtained from the ODE model given by Eq. (19); see the two cyan curves. In the stable regime, these ODE frequencies correspond to the two eigenfrequencies emerging from the Hamiltonian Hopf bifurcation within the PDE model. Note the excellent agreement between these two BdG eigenfrequencies and their ODE counterparts. It is important to note that this agreement is valid for a relatively wide range of $\mu_{b}$ values since the fit of the ODE frequencies, and hence the fit of the parameters $M_{0}$ and $\Lambda$ in Eqs. (10), was performed for each value of $\mu_{b}$. Nonetheless, we checked that, for each chosen value of $\mu_{b}$ (and the corresponding fitted parameters), the PDE and ODE orbits matched for a wide range of initial conditions. This evidences the validity of the ODE model not only qualitatively, but more importantly, quantitatively in its ability to capture the VB evolution within the full PDEs once the parameters of the system are set.
@@
Figure 6: Evolution of an initially centered unstable VB structure for $\mu_{d}=4.25$ and $\mu_{b}\approx 4.38$. The first and second rows of panels depict the (squared root) of the densities for, respectively, the bright and dark components at the times indicated. The blue spiral in the top-left panel corresponds to the orbit followed by the VB structure during its destabilization for $0\leq t\leq 1000$. The third and fourth rows of panels depict, respectively, the phase and vorticity of the dark component. For convenience, the phase panels only show the region inside the Thomas-Fermi radius (outside this region the phases are very complicated as the density rapidly decays to zero). The yellow (black) dots in the vorticity panels depict the location of positively (negatively) charged vortices where the vorticity attains its maximum (minimum) denoted by $w_{M}$ ($w_{m}$). The right panel depicts the evolution of the vorticity by means of isocontours of vorticity in $(x,y,t)$ where the red and green contours correspond to negative and positive vorticities, respectively.
@@
On the other hand, for our chosen parameters and as per our stability analysis, for $\mu_{b}<4.4$ the VB structure centered at the origin is unstable and tends to spiral out. We provide a typical example of the destabilization dynamics in this region in Fig. 6 which depicts the time-evolved solution together with its phase and vorticity profiles. As expected, for initial times ($t\lesssim 1000$), the VB destabilizes from the origin and performs a spiraling-out motion (see overlaid blue curve in the top-left panel) while, approximately, preserving its VB initial shape albeit a small elongation along the spiralling direction. Further down the evolution, this elongation of the VB “patch” becomes more pronounced allowing for the formation of extra vortices inside the VB. Initially, a vortex-anti-vortex pair creation occurs inside the VB complex (see the state at $t=1030$). Then, the corresponding negatively charged vortex keeps trapped within the VB complex while the positively charged vortex, carrying a small amount of bright mass, is ejected out of the VB and runs toward the edge of the condensate orbiting in the opposite direction of the VB motion. Next, as the VB patch continues to stretch, a series of aligned vortices of negative charge are created giving rise to a quasi-1D vortex “patch”. Eventually, the stretched VB patch breaks into two individual VB patches, each carrying a portion of the vortices. We note that, to the best of our knowledge, this intrinsic creation of vortices (i.e., without the need of an external stirrer or potential) is associated with a novel type of VB-creating instability. What seems to be happening here involves the transfer of angular momentum associated with the spiraling motion of the VB into net vorticity since the vortices produced are predominantly of the same charge.
@@
Figure 7: Temporal evolution of the angular momentum for the destabilization dynamics of an initially centered VB depicted in Fig. 6. The initial gradual angular momentum exchange starting around $t\approx 800$ corresponds to the spiralling-out of the VB. As time progresses, the angular momentum of each component tends to vary around the equipartition level $L_{d}\approx L_{b}\approx L_{z}/2$.
@@
In order to garner more information about the emergence of new vortices through the above-mentioned destabilization dynamics of the centered VB complex, we monitor the vorticity in the system. In particular, the angular momentum in each component can be computed using Mukherjee et al. (2020):
@@
where the subscripts denote the first (dark) and second (bright) components. Since the underlying potentials for both components are independent of the angular coordinate, the total angular momentum of the system is conserved:
$L_{z}(t)=L_{d}(t)+L_{b}(t)=L_{z}(0).$ (25)

### v2

@@ I Introduction
The platform of atomic Bose-Einstein condensates (BECs) has been particularly fruitful towards the exploration of nonlinear wave patterns, both solitonic ones, as well as vortical ones involving topological charge Pethick and Smith (2002); Stringari and Pitaevskii (2003). The ability to manipulate external potentials in both space and time Henderson et al. (2009), as well as that of controlling the size and the sign of the atomic scattering length (and hence the coefficient of nonlinear interaction) in time Donley et al. (2001); Pollack et al. (2009) or even in space Kartashov et al. (2011); Di Carli et al. (2020) have paved new directions of unprecedented perturbations and dynamics to the solitary wave patterns that such systems can support.
@@
While these higher-component settings offer considerable additional wealth in terms of the possible states, our aim herein is to revisit two-component settings. There, many solitary waveforms have been considered, ranging from dark-bright to dark-dark and dark-antidark, as well as magnetic solitons Kevrekidis et al. (2015); Kevrekidis and Frantzeskakis (2016); Qu et al. (2016). Not only individual structures of this kind have been experimentally explored, but also small clusters thereof (see, e.g., Ref. Katsimiga et al. (2020)), as well as their collisions Becker et al. (2008). Such two-component settings have recently been leveraged also towards engineering effective attractive interactions in both 1D and 2D, offering the possibility to generate structures such as the Townes soliton Bakkali-Hassani et al. (2021), as well as the Peregrine soliton Romero-Ros et al. (2024).
Here, we focus on the 2D variants of the above solitary waves in the form of vortex-bright (VB) solitons Skryabin (2000); Law et al. (2010). These are sometimes referred to by other names such as baby-skyrmions, filled-core vortices, or massive vortices and have not only been theoretically explored but also experimentally observed Anderson et al. (2000); Leanhardt et al. (2003); Choi et al. (2012). Note that some of these experimental realizations such as the one reported in Ref. Choi et al. (2012) (and corresponding theoretical explorations such as in Ref. Katsimiga et al. (2023)) have probed these states in spinor settings. In recent years there has been a flurry of activity in such systems, seeking to provide a theoretical understanding of the potential dynamics of these structures. Indeed, they have been shown to be able to perform flower-like motions Richaud et al. (2021); Caldara et al. (2023); they have been also explored in generalized radial potentials Wang (2022); Richaud et al. (2022), as well as in the context of rotating traps Richaud et al. (2020) and more recently the case of multiple such states and how they interact has been revisited Richaud et al. (2023) (see also earlier studies in Refs. Pshenichnyuk (2018); Kasamatsu et al. (2016)). It is relevant to point out here that in the context of miscible BECs the interaction of these VB solitons was considered earlier in Ref. Eto et al. (2011). An additional, intriguing in its own right, setting where a superfluid bearing a vortex is coupled to an impurity, i.e., a second component without self-interactions but solely feeling the vortex as an “external potential” was recently proposed in Ref. Edmonds et al. (2021). Both the VB-like state but also higher excited states of the second component impurity were considered therein. Finally, yet another direction touched upon recently in the VB setting is that of higher topological charges, where it was found that the bright component could play a significant role in their dynamical stabilization Patrick et al. (2023).
@@ II.1 Full PDE model
The VB states are studied in the relevant parameter space. Here, the key parameters determining the properties of the VB complex are $\Omega$, $g_{12}$, $\mu_{d}$, and $\mu_{b}$. For example, the powers and the mass ratio depend directly, in a non-trivial manner, on the chemical potentials for a given set of $\Omega$ and $g_{12}$ values.
We find numerically exact stationary states using finite differencing for the spatial discretization together with a Newton-based fixed-point iteration method. First, we find a relevant solution, and then numerical continuation can be deployed along any of the above parameters. The first solution is obtained by embedding a vortex centered at the origin in the first component, we shall discuss the details in Sec. III. We also study the (spectral) stability of these centered VB complexes by the BdG analysis. Then, taking parameter values where the VB complex is stable, we proceed to appropriately seed it away from the trap center and monitor, according to Eq. (1), its dynamical evolution. Our dynamics is carried out using the regular fourth-order Runge-Kutta method. The location of the VB complex is monitored by first computing the respective centers of mass of each component along the $x$ and $y$ directions:
@@
where $\sigma=x,y$. Then, the relative location of the VB with respect to the background cloud of the dark component is recorded as ${\rm CoM}^{\sigma}_{2}-{\rm CoM}^{\sigma}_{1}$.
@@
$-i\left[\begin{array}[]{cccc}\phantom{-}M_{11}&\phantom{-}M_{12}&\phantom{-}M_{13}&\phantom{-}M_{14}\\ -M_{12}^{*}&-M_{11}^{*}&-M_{14}^{*}&-M_{13}^{*}\\ \phantom{-}M_{13}^{*}&\phantom{-}M_{14}&\phantom{-}M_{22}&\phantom{-}M_{34}\\ -M_{14}^{*}&-M_{13}&-M_{34}^{*}&-M_{22}^{*}\\ \end{array}\right]\left[\begin{array}[]{c}a_{1}\\ b_{1}\\ a_{2}\\ b_{2}\end{array}\right]=\nu\left[\begin{array}[]{c}a_{1}\\ b_{1}\\ a_{2}\\ b_{2}\end{array}\right],$
@@
and $M_{12}\equiv\psi_{1}^{2}$, $M_{13}\equiv g_{12}\psi_{1}\psi_{2}^{*}$, $M_{14}\equiv g_{12}\psi_{1}\psi_{2}$, and $M_{34}\equiv\psi_{2}^{2}$. The eigenvector for the corresponding eigenvalue $\nu$ is $[\Psi_{1},\Psi_{2}]^{T}=[(a_{1},b_{1}),(a_{2},b_{2})]^{T}$, and the mode is unstable if Re$(\nu)>0$. As we will see in Sec. III, after stationary solutions are displaced away from the origin and for appropriate parameter regimes, the VB states can exhibit flowering trajectories. To follow potential instabilities emerging from the collision of eigenvalues, for each eigenvector we monitor the sign of the quantity (so-called Krein signature) Kapitula et al. (2004):
@@ II.2 Reduced ODE model for the VB complex
Note that we may also write $L$ in polar coordinates using standard conversion formulas $x^{2}+y^{2}=r^{2}$, $\dot{x}^{2}+\dot{y}^{2}=\dot{r}^{2}+r^{2}\dot{\theta}^{2}$, and $x\dot{y}-y\dot{x}=r^{2}\dot{\theta}$ for $r=r(t)$ and $\theta=\theta(t)$. Comparing Eq. (8) above with the Lagrangian (23) of Ref. Richaud et al. (2021), we find similar structural features with the expression used herein bearing an additional contribution $\left[\mu_{d}-\frac{1}{2}V(r)\right]$ in the second kinetic energy term, while instead of the logarithmic term in Ref. Richaud et al. (2021), we include herein only the quadratic term in its Taylor series. Additionally, their $r$ is normalized as $r/R_{\rm TF}$ for $R_{\rm TF}$ the TF radius of the dark component.
@@
On the other hand, to derive the Hamiltonian formulation of the ODE system (7) we can utilize the Legendre transform $H=\dot{x}p_{x}+\dot{y}p_{y}-L$ with $p_{\sigma}=\frac{\partial L}{\partial\dot{\sigma}}$ for $\sigma=x,y$, to obtain
$H$ $=$ $\frac{1}{2M_{0}}\left(p_{x}^{2}+p_{y}^{2}-(yp_{x}-xp_{y})\left(\mu_{d}-\frac{1}{2}V(r)\right)\right)$
$+\frac{r^{2}}{8M_{0}}\left(\mu_{d}-\frac{1}{2}V(r)\right)^{2}-\frac{\Lambda}{4}r^{2}.$
@@
$V_{\rm eff}(r)=\frac{1}{2}M_{0}r^{2}\dot{\theta}^{2}-\frac{\Lambda}{4}r^{2}.$
@@
$M_{0}(\ddot{r}-r\dot{\theta}^{2})=-\rho(r)r\dot{\theta}+\frac{\Lambda}{2}r,$
@@
The smaller frequency $\omega_{\rm pr}\equiv\dot{\theta}_{-}$ corresponds to the expected precessional frequency of a circle, and reduces to the expected single-component precession frequency asymptotically in the limit of $M_{0}\rightarrow 0$. Here, given the presence of the second component, we also obtain a larger frequency $\dot{\theta}_{+}$ that also pertains to a precessional motion, but indeed one whose frequency diverges in the $M_{0}\rightarrow 0$ limit.
@@
$\dot{y}(0)=\frac{1}{2M_{0}}\left(\rho(r_{0})x_{0}\pm\sqrt{\rho(r_{0})^{2}x_{0}^{2}-2M_{0}\Lambda r_{0}^{2}}\right).$
@@
for radial and precessional frequencies $(\omega_{r},\omega_{\rm pr})$ and $R$ and $A_{0}$ being small perturbations off of the circular motion. Plugging Eqs. (LABEL:eq:estRTheta) into the first order system above gives
@@
In the next section, we will make the correspondence between the PDE model and its effective ODE reduction using the following program. Since we expect frequencies $\dot{\theta}$ from the ODE model to correspond to purely imaginary frequencies in the PDE model, we use the BdG bifurcating eigenvalues $\nu$ and the correspondence $\dot{\theta}={\rm Im}(\nu)$ in order to estimate the ODE parameter value of $\Lambda$ in terms of $M_{0}$. The value of $M_{0}$ is then determined by minimizing the sum of squares of differences between the PDE and ODE trajectories. For a given pair of $M_{0}$ and $\Lambda$ one then finds initial velocities of the ODE model that best align to a given PDE trajectory. In general, a larger initial velocity gives flowering trajectories with more looping on the petals so this can be used to guide best choices of these initial velocities in the ODE. For the obtained flowering trajectories, we will also show examples comparing the ODE trajectories according to Eq. (7) versus the results of the linearized predictions in Eqs. (LABEL:eq:estRTheta) and (LABEL:eq:omegaA).
@@ III.1 Steady states and stability for the PDE model
Let us now track the BdG eigenvalue spectrum for the centered VB structure as a function of the parameter $\mu_{b}$. Figure 5 depicts a typical spectrum for $g_{12}=1.6$ and $\mu_{d}=4.25$. Note that Re$(\nu)>0$ indicates an instability. Interestingly, as the spectrum shows, there is a Hamiltonian Hopf bifurcation as $\mu_{b}$ decreases. For $\mu_{d}=4.25$ and values $g_{12}=1.4,1.6,1.75$, the bifurcations occur, respectively, at $\mu_{b}\approx 4.349,4.383,4.403$ and where according to Fig. 2 the respective critical mass ratios are $M^{*}\approx 0.057,0.051,0.048$. This (reverse) Hamiltonian Hopf bifurcation evidences that the lowest two oscillatory (purely imaginary) eigenfrequencies collide as $\mu_{b}$ decreases giving rise to an unstable (complex) quartet. The presence of a complex quartet instability indicates that, in this parametric regime, the VB will destabilize from the center of the trap along an spiraling-out trajectory as we will see below soon. According to the color-coded eigenvalues in the bottom panel of Fig. 5, the region of instability lies where the lower branch is colored black for $\mu_{b}\lesssim 4.383$ and where the mass ratio $M>M^{*}$ exceeds the critical value $M^{*}\approx 0.051$. However, it is interesting that, with the parameters provided in Fig. 5, spiraling-out is also present for $\mu_{b}$ values as high as $\mu_{b}<4.4$ when the VB structure is displaced by a large value such as $x_{0}=2.5$. This tends to suggest that precessional circular orbits may also be prone to instabilities. This is a topic of particular interest in its own right, meriting an examination of such states in a co-rotating frame (as was done, e.g., in the work of Ref. Pelinovsky and Kevrekidis (2012)), yet this type of distinct study is an interesting topic for future work.
@@
Focusing on initially centered VB solutions, according to our BdG spectral analysis, starting infinitesimally close to the center, spiral-out trajectories will only be present in the $\mu_{b}$-interval where the stationary solutions are unstable. For instance, the VB dynamics for an initial condition $(x_{0},y_{0})=(2.5,0)$ in the stable regime ($\mu_{b}\approx 4.4$) displays flowering trajectories that have roughly two petals and as $\mu_{b}$ increases the number of petals increase. In Fig. 5, we overlay the two frequencies obtained from the ODE model given by Eq. (13) (see the two cyan curves). In the stable regime, these ODE frequencies correspond to the two eigenfrequencies emerging from the Hamiltonian Hopf bifurcation within the PDE model. Note the excellent agreement between these two BdG eigenfrequencies and their ODE counterparts. It is important to note that this agreement is valid for a relatively wide range of $\mu_{b}$ values since the fit of the ODE frequencies, and hence the fit of the parameters $M_{0}$ and $\Lambda$ in Eqs. (7), was performed for each value of $\mu_{b}$. Nonetheless, we checked that, for each chosen value of $\mu_{b}$ (and the corresponding fitted parameters), the PDE and ODE orbits matched for a wide range of initial conditions. This evidences the validity of the ODE model not only qualitatively, but more importantly, quantitatively in its ability to capture the VB evolution within the full PDEs once the parameters of the system are set.
@@
Figure 6: Evolution of an initially centered unstable VB structure for $\mu_{d}=4.25$ and $\mu_{b}\approx 4.38$. The first and second rows of panels depict the (square root) of the densities for, respectively, the bright and dark components at the times indicated. The blue spiral in the top-left panel corresponds to the orbit followed by the VB structure during its destabilization for $0\leq t\leq 1000$. The third and fourth rows of panels depict, respectively, the phase and vorticity of the dark component. For convenience, the phase panels only show the region inside the Thomas-Fermi radius (outside this region the phases are very complicated as the density rapidly decays to zero). The yellow (black) dots in the vorticity panels depict the location of positively (negatively) charged vortices where the vorticity attains its maximum (minimum) denoted by $w_{M}$ ($w_{m}$). The right panel depicts the evolution of the vorticity by means of isocontours of vorticity in $(x,y,t)$ where the red and green contours correspond to negative and positive vorticities, respectively.
@@
On the other hand, for our chosen parameters and as per our stability analysis, for $\mu_{b}<4.4$ the VB structure centered at the origin is unstable and tends to spiral out. We provide a typical example of the destabilization dynamics in this region in Fig. 6 which depicts the time-evolved solution together with its phase and vorticity profiles. As expected, for initial times ($t\lesssim 1000$), the VB destabilizes from the origin and performs a spiraling-out motion (see overlaid blue curve in the top-left panel) while, approximately, preserving its VB initial shape albeit a small elongation along the spiraling direction. Further down the evolution, this elongation of the VB “patch” becomes more pronounced allowing for the formation of extra vortices inside the dark component of the VB. Initially, a vortex-anti-vortex pair creation occurs inside the VB complex (see the state at $t=1030$). Then, the corresponding negatively charged vortex keeps trapped within the VB complex while the positively charged vortex, carrying a small amount of bright mass, is ejected out of the VB and runs toward the edge of the condensate orbiting in the opposite direction of the VB motion. Next, as the VB patch continues to stretch, a series of aligned vortices of negative charge are created giving rise to a quasi-1D vortex “patch”. Eventually, the stretched VB patch breaks into two smaller VB patches, each carrying a portion of the vortices. We note that, to the best of our knowledge, this intrinsic creation of vortices (i.e., without the need of an external stirrer or potential) is associated with a novel type of VB-creating instability. What seems to be happening here involves the transfer of angular momentum associated with the spiraling motion of the VB into net vorticity since the vortices produced are predominantly of the same charge. This is partially reminiscent of the process of generation of persistent currents (see, e.g., Ref. Law et al. (2014)). A slightly different in spirit, yet also intriguing, recent example of vortical phase structure produced in the bright components of VB complexes in rotating traps, can be found in Ref.Chaika et al. (2023). The latter so-called ghost vortices can be rendered real through suitable interaction quench protocols as discussed in Ref. Chaika et al. (2023).
@@
Figure 7: Temporal evolution of the angular momentum for the destabilization dynamics of an initially centered VB depicted in Fig. 6. The initial gradual angular momentum exchange starting around $t\approx 800$ corresponds to the spiraling-out of the VB. As time progresses, the angular momentum of each component tends to vary around the equipartition level $L_{d}\approx L_{b}\approx L_{z}/2$.
@@
In order to garner more information about the emergence of new vortices through the above-mentioned destabilization dynamics of the centered VB complex, we monitor the vorticity in the system. In particular, the total angular momentum in each component can be computed using Mukherjee et al. (2020):
@@
where the subscripts denote the dark and bright components. Since the underlying potentials for both components are independent of the angular coordinate, the total angular momentum of the system is conserved:
$L_{z}(t)=L_{d}(t)+L_{b}(t)=L_{z}(0).$